OpenStudy (anonymous):

If the circumference of a circle is 300 ft, what is the radius?

OpenStudy (anonymous):

\[C=2(\pi)r\]

OpenStudy (anonymous):

so, \[300=2(\pi)\]

OpenStudy (anonymous):

**2(pi)r sorry, messed up

OpenStudy (anonymous):

300= 2(pi)r 300/2 = (pi)r 150 = (pi)r 150/(pi) = r

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

approx. 47.746

OpenStudy (anonymous):

No problem.